in a bold attempt to test the market, Kate raises her price of hot dogs from R1 to R1.25 and her sales fall from 150days to 125days. From the data given what is the point of elasticity of demand for hot dogs at the new price? Assume the demand curve is linear


1
Expert's answer
2021-06-14T23:49:01-0400

Percentage change in price = (1.25×100)÷1= 125%


Percentage change in demand= (125×100)÷150 = 83.3%


Point of elasticity of demand= (percentage of demand change) ÷(percentage of price change)


= (100-83.3)÷(100-125)

 (16.7)÷(-25)

-0.668%